OpenStudy (anonymous):

help me I cant figure this out! f'(x) = 3x^2 - 6x + 2. The point (1,11) lies on y = f(x).

OpenStudy (xapproachesinfinity):

what do you want to do

OpenStudy (loser66):

question?

OpenStudy (xapproachesinfinity):

I'm assuming you want to find f(x) given f'(x)

OpenStudy (jhannybean):

So you're given the equation of a tangent line and a point that lies on the curve where that tangent line may intersect?

OpenStudy (xapproachesinfinity):

that's not really a tangent line, it is slope of f at any given x

OpenStudy (xapproachesinfinity):

at this stage we cannot help you out with anything if you don't write your question

OpenStudy (xapproachesinfinity):

do we help to write your question? lol

OpenStudy (loser66):

f'(x) = 3x^2 -6x+2 \(\color{red}{Does}\) the point (1,11) lie on y = f(x)?

OpenStudy (mathmate):

Hints: dy/dx=f'(x)=3x^2-6x+2 \(\int f'(x)dx=f(x)+C\) use (1,11) to evaluate C.